OpenStudy (anonymous):

4. k2 – 81h2 (1 point) A (k – 9h2)(k + 9) B (k + 9h)(k + 9h) C h2(k + 9)(k – 9) D (k + 9h)(k – 9h)

OpenStudy (anonymous):

its a problem that would involve the application of the difference of two square that says that: a^2-b^2=(a-b)(a+b). from your question the next step will be to factorise into the form a^2 - b^2...so we will have: (k)^2 -(9h)^2=(k+9h)(k-9h)

OpenStudy (anonymous):

Thanks so much!